OpenStudy (anonymous):

Okay, so some somehow I am just completely stuck this question. This is one of the examples from my textbook that I just can't solve and would probably help me solve the others I'm stuck on. Solving by Substitution 3x+4y=12 2x+y=10 I keep getting this: 3x+4(-2x+10)=12 3x-8x+40 -5x+40=12 -5x=-32 x=6.4 Am I missing something staring me in the face? Because the answer should be x=5.6

OpenStudy (kewlgeek555):

Are you studying slopes and linear functions and stuff?

OpenStudy (kewlgeek555):

Because to me that looks lie the standard form of a linear function, which is Ax - By = C.

OpenStudy (anonymous):

Well, it's in the Linear Systems Chapter under Solving Systems Algebraically.

OpenStudy (anonymous):

I should note, this is Algebra 2 if that helps.

OpenStudy (kewlgeek555):

Okay then. I think what they mean is to find x first then y. This is how you do it!:\[3x+4y=12\]\[3x+4(0)=12\]\[3x=12\]\[12\div3=x\]\[4=x\]Now, you are not finished yet. All you have to do next is put it in a coordinate plane thingie: (4, 0). Now - did you see what I did with the y? I just made it zero. Do the same thing with x, when finding y. Try to do it yourself! :-)

OpenStudy (kewlgeek555):

And OMG, Algebra II. I am in Algebra I, but I just finished a chapter called "Linear Functions". But I guess it would also be linear systems because I also learned linear models, linear equations, etc...

OpenStudy (kewlgeek555):

Please tell me when you are done.

OpenStudy (kewlgeek555):

...Or if you need assistance.

OpenStudy (kewlgeek555):

I'll tell you what to do after you solve for y.

OpenStudy (anonymous):

Doing it on some paper, just a sec.

OpenStudy (anonymous):

Okay, well I see what you're saying but this is a textbook example, and it is telling me that x=5.6, and somehow I've manged to get x=6.4 and x=-1.6 the two times I solved it. Haven't even gotten to y yet. That is the whole issue.

OpenStudy (kewlgeek555):

Hm. So x HAS to equal 5.6? Well, if x = 5.6, can't you plug in 5.6 and then solve for y? Or is 5.6 the answer?

OpenStudy (anonymous):

5.6 is x which I then need to use to find y, I think I see where I went wrong so I'm going to try again real quick.

OpenStudy (kewlgeek555):

Okay. :-)

OpenStudy (anonymous):

Figured out where I went wrong, 12-40=-28, not 32. Makes a big difference and 4*10 does not =4. Moment of stupidity. Got it now, thanks for the help lol.

OpenStudy (kewlgeek555):

Okay, you are welcome. :-D